OpenStudy (anonymous):

The length of a rectangle is twice the width. The area is 162^2yd. Find the length and the width

OpenStudy (anonymous):

I think: 2w^2=162^2 w^2=162*81 w=81sqrt(2) l=162sqrt(2)

OpenStudy (anonymous):

your answer is correct CrazyEinstein.

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

CrazyEinstein and renanmit's conclusions are incorrect. { L = 2*w, L*w = 162 } L = 18 and w = 9 L * w = 162 The length times the width using CrazyEinstein's answer is:\[162 \sqrt{2}*81 \sqrt{2}=26244 \]

OpenStudy (stormfire1):

A = L*W L = 2W A = 2W * W A = 2W^2 A/2 = W^2 81 = W^2 W = 9 and since the length is 2W, the length will be 18. So the final equation would look like 162 = 18 * 9

OpenStudy (anonymous):

162^2yd means 162^2 or 162 ? if it is 162^2, mine is right. if it is 162, stormfire1's is right.

OpenStudy (stormfire1):

Since it's an area, it's not 162^2....it's 162 square yards

OpenStudy (stormfire1):

That's how I read it anyway

OpenStudy (anonymous):

Here is how WolframAlpha.com reads l=162sqrt(2) http://www.wolframalpha.com/input/?i=l%3D162sqrt%282%29

OpenStudy (stormfire1):

I'd bet a lot of $ on the fact that she meant \[162 yd^2\]:)

OpenStudy (anonymous):

I agree. That is how I interpreted the given area. The go around regarding l=162sqrt(2) refers to @ CrazyEinstein's answer for the length.

OpenStudy (anonymous):

Haha, there is no need to bet. it is just a matter of clarification.